Determine the speed of the runner completing the 100 m loop twice in a total time of 50 seconds?

a)  0.4 m/s b) 0.4 m/s2 c. 4 m/s d. 4 m/s2




1
Expert's answer
2021-03-29T09:06:46-0400
"v=\dfrac{s}{t}=\dfrac{2\cdot100\ m}{50\ s}=4.0\ \dfrac{m}{s}."

Answer:

 c. 4 m/s